A country’s labor force is the sum of the number of employed and unemployed workers. The accompanying table provides data on the size of the labor force and the number of unemployed workers for different regions of the United States.





Labor force

Unemployed

Region

April 2013

April 2014

April 2013

April 2014

Northeast

28,407,200

28,288,900

2,174,400

1,781,300

South

56,787,800

57,016,400

4,089,900

3,363,800

Midwest

34,320,000

34,467,000

2,473,700

2,109,000

West

36,122,200

36,307,300

2,940,800

2,535,700



Calculate the number of workers employed in each of the regions in April 2014.


Number of workers employed in Northeast region in April 2014

Number of workers employed in South region in April 2014

Number of workers employed in Midwest region in April 2014

Number of workers employed in West region in April 2014
1
Expert's answer
2021-08-29T16:31:26-0400

Number of workers employed in Northeast region in April 2014

  • The number of employed workers = Labor force – the number of unemployed workers.
  • But,
  • Labor force = 28,288,900
  • The number of unemployed workers = 1,781,300
  • Hence,
  • The number of employed workers = 28,288,900 – 1,781,300 = 26,507,600.

Number of workers employed in South region in April 2014

  • The number of employed workers = Labor force – the number of unemployed workers.
  • But,
  • Labor force = 57,016,400
  • The number of unemployed workers = 3,363,800
  • Hence,
  • Number of employed workers = 28,288,900 – 1,781,300 = 53,652,600.

Number of workers employed in Midwest region in April 2014

  • The number of employed workers = Labor force – the number of unemployed workers.
  • But,
  • Labor force = 34,467,000
  • The number of unemployed workers = 3,363,800
  • Hence,
  • Number of employed workers = 34,467,000– 2,109,000 = 32,358,000.

 

Number of workers employed in West region in April 2014

  • The number of employed workers = Labor force – the number of unemployed workers.
  • But,
  • Labor force = 36,307,300
  • The number of unemployed workers = 2,535,700
  • Hence,
  • Number of employed workers = 36,307,300 – 2,535,700 = 33,771,600.